suppose you roll two dice find the probability of not rolling a sum of 7
Answer by ewatrrr(24785)


Hi
suppose you roll two dice find the probability of not rolling a sum of 7
1 - P(sum of 7) = 1-6/36 = 30/36 = 5/6
___1_2_3__4__5__6
1|_2_3_4__5__6__7
2|_3_4_5__6__7__8
3|_4_5_6__7__8__9
4|_5_6_7__8__9_10
5|_6_7_8__9_10_11
6|_7_8_9_10_11_12